OpenStudy (anonymous):

y^2 = (x-1)/(x+1). i need to find dy/dx

OpenStudy (anonymous):

i guess you could take the square root, and then find the derivative of that ugly mess. or else you could write \[2yy'=\frac{d}{dx}\left[\frac{x-1}{x+1}\right]\] take the derivative, and divide by \(2y\)